### v1

# First Heuristic Then Rational: Dynamic Use of Heuristics in Language Model Reasoning

Abstract
Multi-step reasoning is widely adopted in the community to explore the better performance of language models (LMs). We report on the systematic strategy that LMs use in this process. Our controlled experiments reveal that LMs rely more heavily on heuristics, such as lexical overlap, in the earlier stages of reasoning when more steps are required to reach an answer. Conversely, as LMs progress closer to the final answer, their reliance on heuristics decreases. This suggests that LMs track only a limited number of future steps and dynamically combine heuristic strategies with logical ones in tasks involving multi-step reasoning. [^1] The code/data will be made public upon acceptance.
1 Introduction
When facing complex tasks, humans tend to seek shallow, heuristic solutions first; and, once these attempts are revealed to fail or elicit another reasonable solution, they switch to being more rational Erickson and Mattson (1981); Frederick (2005). This systematic behavior helps us to predict how humans will tackle new problems. Given such a view, when it comes to predicting the behavior of language models (LMs) Madaan and Yazdanbakhsh (2022); Ye et al. (2023), the following question naturally arises—Do LMs also use a similar systematic strategy to solve complex tasks, or is their strategy totally different from humans, or do they have no such strategies? This study explores an answer to this question. Investigating LMs’ strategic behavior in problem-solving is expected to provide a new perspective on LMs’ reasoning mechanism. It may also address general concerns that current neural models tend to overly rely on superficial, heuristic cues and may end up with irrational conclusions Du et al. (2022); Lai et al. (2021); Jia and Liang (2017); Ye et al. (2023); Chen et al. (2024).
Specifically, in the course of step-by-step reasoning, they tend to rely on shallow, short-sighted heuristic preferences in choosing rules in the early phase and dynamically switch to be more rational and goal-oriented to make the right choice to reach the goal. This highlights a severe limitation of present LMs, including GPT-4 OpenAI (2023), in searching for a solution to multi-hop reasoning tasks, particularly when tasks require many-step long solutions.
Figure 1: Illustration of the systematic strategy we discovered in language models (LMs). When the goal is distant from the current state in a multi-step reasoning process, the models tend to rely on heuristics, such as superficial overlap, which can lead them in the wrong direction. In contrast, when the goal is within a limited distance, the models are more likely to take rational actions to reach the goal.
Figure 2: Overview of the task setting. Given premises and a question, a model answers the question step-by-step (left part). Through each reasoning step $t$ of selecting/paraphrasing relevant premise $p_{k}\in P$, the available facts $\bm{z}$ are enriched, and if that step is necessary to reach the answer, the distance to the answer $d$ decreases (right part).
2 Task
As a controlled testbed to analyze LM’s reasoning ability, we adopt an arithmetic reasoning task (Figure 2 left). We will use both natural and artificially controlled datasets in the experiments, but let us use the latter, more formal examples to explain the task overview.
Arithmetic reasoning task:
The problem consists of a set of premises $P=\{p_{1},\cdots,p_{k}\}$ and a question $q$. Each premise describes either type of fact: (i) Person A has $n$ items (A=$n$), or (ii) Person B has $n$ more/less items than A has (B=A+$n$ or B=A-$n$). The question asks the exact number of certain items a particular person ultimately has (How many items does B have?). Here, one should consider multiple premises to derive the final answer, e.g., A=3;B=2+A;B=2+3=5. Notably, some premises are irrelevant to the answer; thus, models have to track which premise is necessary to reach the final answer.
Reasoning step:
Let $f$ be a model that is instructed to solve the task step-by-step. In each reasoning step $t$, the model $f$ selects a particular premise $p_{i}\in P$ and paraphrases it into a new fact $z_{t}$ by eagerly resolving reference expressions based on the already stated facts $\bm{z}_{<t}=[z_{1},\cdots,z_{t-1}]$:
$f(P,q,\bm{z}_{<t})=(p_{i},z_{t})\;\;\,\mathrm{.}$ (1)
For example, in Figure 2, when $p_{2}$ “Walter has 2 more apples than Peggy.” is selected at a particular reasoning step, the respective $z_{t}$ should be “Walter has 2+5=7 apples.” if $\bm{z}_{<t}$ already contains the number of apples Peggy has, i.e., $p_{1}$. [^2] If the reference can not be resolved with $\bm{z}_{<t}$, the model just repeats the selected premise $p_{i}$ as $z_{t}$. Starting with an empty set of stated facts $\bm{z}=\{\}$, the model recursively performs a reasoning step and can stop when outputting a special symbol EOS or providing an answer to the question $q$. Here, we denote the whole history of selected premises as $\bm{h}=[p_{i},\cdots,p_{j}]\in P^{*}$, where $P^{*}$ is Kleene closure of $P$. Its $t$-th element $h_{t}$ is the premise employed to derive the $t$-th reasoning step. Henceforth, we call $\bm{h}$ reasoning steps and focus on the ability to search the right $\bm{h}$.
Solutions:
Each problem has a set of solutions $H^{\circ}\subset P^{*}$. Specifically, the final stated fact $z_{-1}$ generated from the final reasoning step $h_{-1}$ of the solution $\bm{h}\in H^{\circ}$ should provide a concrete number asked by the question $q$. Figure 2 illustrates such a set of solutions $H^{\circ}$ as the steps leading to the final states of the state transition graph (right part of Figure 2).
Minimal solution:
Within the set of solutions, there is only one minimal solution $\bm{h}^{*}\in H^{\circ}\subset P^{*}$. To define $\bm{h}^{*}$, let us first introduce a distance to the answer. In each reasoning step $t$, one can determine the minimum number of remaining reasoning steps to reach the answer $d\in\mathbb{N}$, given $\bm{h}_{\leq t}\in P^{*}$ and the initially provided premises $P$. Intuitively, $d$ can be derived from a state transition graph and the minimum number of transitions to the closest final states, as shown in Figure 2 (right part). Here, we denote the mapping function from $\bm{h}_{\leq t}$ to $d$ as $g:P^{*}\rightarrow\mathbb{N}$. For example, $g([p_{2},p_{1},p_{2}])=1$ in Figure 2. Minimal solution $\bm{h}^{*}$ satisfies $\forall t\ g(\bm{h}^{*}_{\leq t})<g(\bm{h}^{*}_{\leq t-1})$; that is, $\bm{h}^{*}$ does not contain any irrelevant step to approach the answer. For example, the MS reasoning steps in Figure 2 are $[p_{1},p_{2},p_{4}]=\bm{h}^{*}$.
Targeted ability of LMs:
We evaluate LMs’ ability to derive the minimal solution $\bm{h}^{*}$ as instructed by 4-shot examples 10. Notably, we do not care about the ability to correctly introduce a new fact $z_{t}$ (Eq. 1), e.g., the accuracy of arithmetic operation (e.g., 5+2=7), but separately focus on their search strategy to select the relevant premise to perform the next reasoning step.
3 Heuristics
Given existing studies on LMs’ use of heuristics (§ 5), we focus on the following types of heuristics:
Lexical overlap between premise and question (Overlap):
Whether models tend to select premises with the same person name (PN) as the one in question is targeted. For example, given a question “how many apples Judy has,” premises such as “Judy’s mother got 3 apples” can be a distractor when it is unnecessary to reach the answer.
Position of premise (Position):
Whether models tend to select the premise in the initial position of context is targeted.
Grammatical feature of premise (Neg.):
Given that a specific grammatical feature, e.g., whether the sentence has a negation word, is often a superficial cue Du et al. (2021); Niven and Kao (2019), we specifically analyze the bias that models avoid selecting premise with negation.
We append distractor premises that are irrelevant to answering the question but presumably preferred by the model w.r.t. some heuristics to the input $P$. We analyze whether LMs select such distractors to understand their search strategy (§ 4).
4 Experiments
We use four representative variants of LLMs: text-bison-001 version of Google’sPaLM2 Anil et al. (2023), Llama2-13B Touvron et al. (2023), gpt-3.5-turbo-0125 and gpt-4-0613 snapshots of OpenAI’s GPT-3.5-turbo OpenAI (2022) and GPT-4 OpenAI (2023).
4.1 Preliminary experiments
First, we confirm that LLMs exploit specific heuristics during step-by-step reasoning in natural and artificially-controlled datasets.
Settings:
We use two datasets: GSM8K Cobbe et al. (2021) (App. A) and artificially-controlled data with 4-step arithmetic reasoning (App. B). Each instance in either dataset has at least one premise $\tilde{p}\not\in\bm{h}^{*}$ that is irrelevant to inducing the answer (i.e., distractor).
We randomly select one such irrelevant premise (Base), [^3] Strictly speaking, we randomly selected the premise that does not match any of the three heuristics. modify it to match a particular heuristics (e.g., move it to the first position), and observe whether such a modification makes it more attractive for LLMs to select during step-by-step reasoning. We report the percentage of the problems, where the model selected the distractor at least once during reasoning. We separately run the experiments for each type of heuristic rather than adding all types of distractors to the input at once.
GSM8K Artificial data
Models Base Over. $\uparrow$ Pos. $\uparrow$ Neg. $\downarrow$ Base Over. $\uparrow$ Pos. $\uparrow$ Neg. $\downarrow$
PaLM2 18.4% 57.9% 19.7% 17.1% 10.3% 42.3% 12.0% 4.3%
Llama2 43.2% 69.7% 50.0% 14.5% 32.6% 67.7% 33.0% 41.7%
GPT-3.5 35.5% 67.1% 36.8% 22.4% 21.0% 15.0% 49.0% 0.0%
GPT-4 21.1% 35.5% 22.4% 21.1% 0.0% 0.01% 0.0% 0.0%
Table 1: The percentage of the problems where the model selected a distractor $\tilde{p}$ in step-by-step reasoning. Over. and Pos. denote the Overlap and Position biases.
Results:
Table 1 shows the results. The Base results (baseline) indicate how many times the randomly sampled distractor $\tilde{p}$ is selected during reasoning. The scores for the distractors with heuristic features (Over., Pos., Neg. in Table 1) are generally higher (lower for Neg.) than the Base scores across models and datasets. This indicates that LMs, on average, tend to rely on our targeted heuristics (§ 3). Note that, interestingly, different models yield somewhat different preference towards distractor types; for example, Llama2 and GPT-3.5 are more biased premise position than PaLM2 and GPT-4.
4.2 Main experiments
Then, we further investigate the strategies of LMs to employ heuristics. We hypothesized that the more distant the reasoning step is from the answer (higher $d$), the more heavily models rely on heuristics. This is motivated by the larger gap between questions and available knowledge, leading to difficulty. Thus, models have to rely on more primitive and heuristic factors.
Context: Peggy has 5 apples. Walter has 2 more apples than Peggy. Judy’s mother has 3 less apples than Peggy. Judy has 4 more apples than Walter has.
Question: How many apples does Judy have?
Table 2: Example of a distractor examined in § 4.2. Suppose that $h^{*}_{1}$ is “Peggy has 5 apples.” One of the two candidates with “Peggy” is the correct next step $h^{*}_{2}$ (green), and the other is a distractor (orange).
Figure 3: How frequently a particular distractor is selected (y-axis: $r$) in each reasoning step (x-axis: $d$).
Distractor and evaluation:
To identify in which steps heuristics are more likely to be exploited, ideally, one should design a distractor equally attractive to all the reasoning steps in $\bm{h}^{*}$ and analyze when it is selected during the reasoning to facilitate a fair step-wise comparison; however, such a distractor is inherently difficult to implement. Instead, we add multiple distractors $\tilde{P}$ to the artificial data; each of them $\tilde{p}_{t}\in\tilde{P}$ correspond to each reasoning step $h^{*}_{t}$ in the sense that both share the same person name that appeared in the previous step $h^{*}_{t-1}$ (Table 2). [^4] To rule out the shortcut cue regarding the reference frequencies of each person name, we further added distractor premises to make the frequencies uniform. Similar to § 4.1, we further modify each distractor $\tilde{p}_{t}\in\tilde{P}$ to match each heuristic (Overlap with question or Position in § 3). [^5] We excluded the Neg. (avoidance) bias because if a model avoids negation in the latter step, we could not distinguish whether it was due to heuristic or rational search. In evaluation, for each $t$, partial correct reasoning steps $\bm{h}^{*}_{<t}$ are teacher-forced to a model, and we analyze whether the model selects the right next step $h^{*}_{t}$ or its respective distractor $\tilde{p}_{t}$. We calculate the frequency #()#\#($\cdot$)# ( ⋅ ) of models’ selecting $\tilde{p}_{t}$ and $h^{*}_{t}$; then, their ratio $r=\frac{\#\tilde{p}_{t}}{\#h^{*}_{t}}$ is reported. The chance rate (i.e., random premise selection) is 0.5. Our hypothesis is that the more current step $t$ is distant from the goal (i.e., the larger $d=g(\bm{h}^{*}_{t})$), the more frequently the distractor is selected (i.e., the higher $r$).
Data:
We used 5-step artificial reasoning data and excluded the first ($d=5\rightarrow 4$) and the last ($d=1\rightarrow 0$) steps from evaluation regarding their special properties. GSM8K was excluded due to the infeasibility of controlled distractors.
Results:
The results are shown in Figure 3. The x-axis is the remaining steps $d$ to the goal, and the y-axis is the ratio $r$. The more distant the current step is from the answer (larger $d$), the more frequently the distractor is selected (larger $r$) which is typically above the chance rate. PaLM2 and GPT-4 exhibited particularly clear tendencies of the negative slopes between $d$ and $r$. These suggest the systematic behavior that LMs tend to rely more on heuristics in the earlier reasoning steps.
5 Related work
Multi-step symbolic reasoning:
Given the historical goal of neuro-symbolic integration Hamilton et al. (2022) and the increasing relevance of neural LMs Fang et al. (2024), whether neural models can emulate particular symbolic operations (e.g., search over the graph) has been a key question Yao et al. (2023). In contrast to delineating what symbolic tasks are (im)possible for LMs by varying task complexities Clark et al. (2020), we investigate the inherent, systematic biases in solving a certain reasoning task.
Heuristics in LM:
Neural models have typically been distracted by superficial biases Du et al. (2022). For example, they tend to use superficial features such as overlaps Lai et al. (2021); Sen and Saffari (2020), positional Ko et al. (2020), and specific syntactic category features Du et al. (2021); Niven and Kao (2019) even with chain-of-thought prompting Madaan and Yazdanbakhsh (2022); these motivated our experimental settings.
Search algorithm:
Finding the shortest path between the start and the goal on a graph is a standard problem in computer science Russell and Norvig (2016). Our investigation of LMs on the arithmetic tasks can be seen as characterizing LMs’ biases as a search algorithm. The use of heuristics in graph search is, more or less, related to the A* search algorithm Hart et al. (1968), although heuristics in A* search is a more narrow concept regarding the distance to the goal than those employed by LMs, e.g., position bias.
6 Conclusion
We have found a systematic strategy for the use of heuristics in LMs’ multi-step reasoning—a dynamic transition from a heuristic to a rational reasoning strategy during the course of LMs’ step-by-step reasoning. These results are hopefully helpful for researchers to understand their underlying mechanism as well as for LM users to consider the inherent biases systems have.
Limitations
This study focused only on four specific language models and two arithmetic tasks. Increasing the coverage of models and tasks are obviously possible future direction, although we ensured that our finding generalizes at least several models and task settings. Nevertheless, in particular, Section 4.2 used only artificial datasets for fair comparison. Constructing a controlled, but natural dataset to evaluate the reasoning strategies of LMs should be encouraged. Furthermore, our findings are based solely on the model’s output texts. Elucidating the underlying mechanisms and the source of these behaviors (e.g., training data) should be investigated in the future work.
Ethics statement
This paper does not involve ethical concerns in the sense that we (i) did not conduct human experiments, (ii) just created artificial data without any potentially harmful contents, and (iii) did not address tasks related to ethically sensitive topics (i.e., arithmetic reasoning).
Acknowledgements
We thank for the member of Tohoku NLP Group for their cooperation in this research.
This work was supported by JSPS KAKENHI Grant Numbers 21K21343 and 24K16077, JST CREST Grant Number JPMJCR20D2 and JST SPRING Grant Number JPMJSP2114.
Appendix A Experimental setting with GSM8K (§ 4.1)
A.1 Dataset construction process
As described in 4.1, we modify the existing multi-hop numerical reasoning dataset, GSM8K (distributed under the MIT license), to construct the evaluation dataset. The dataset construction process is divided into two steps: 1. Extracting samples that can be used as evaluation data for this study, and 2. Inserting distractors according to the heuristic we want to evaluate for each extracted problem statement.
A.1.1 Sample extraction
Context: Jamesname decides to run 3num sprints 3num times a week. Hepronoun runs 60num meters each sprint.
Question: How many total meters does hepronoun run a week?
Person’s Names: James
Numbers: 3,60
Table 3: Extraction of names, personal pronouns, and numbers on GSM8K.
We extract samples from the GSM8K evaluation dataset for this study following the process below:
  1. 1.
    We manually create a list of 50 person names (PNs) from a subset of the GSM8K evaluation dataset.
  2. 2.
    Using regular expressions, we identify PNs from this list, as well as pronouns and numerical expressions present in each sample.
  3. 3.
    We extracted samples that included exactly one from our list in both the context and the question, and where either the PN or a pronoun appears in the question (e.g., Table 3).
  4. 4.
    We replaced all pronouns within the extracted samples with the corresponding person’s name.
A.1.2 Distractor insertion
Subsequently, we added distractors to the extracted samples according to each heuristic, thereby constructing a total of 76 samples for the evaluation dataset. Below, we will describe the process of creating the evaluation dataset for each heuristic.
Base
As a baseline, we insert a template-based random distractor (i.e., $\tilde{p}$) into each sample. The distractor was created using the following steps:
  1. 1.
    We randomly selected one sentence from the sample that included both a PN or pronoun and a numerical expression.
  2. 2.
    We replaced the PN or pronoun in the selected sentence with a placeholder, [name].
  3. 3.
    We replaced the numerical expression in the selected sentence with a placeholder, [num].
  4. 4.
    We replaced [name]  with a randomly selected name from the list of PNs created in A.1.1, excluding the name already present in the sample.
  5. 5.
    We replaced [num]  with another value. [^6] The replacement number was calculated by multiplying each number appearing in the sentence by either 0.5, 0.8, 1.2, 1.5, or 2 and then rounding down to the nearest whole number.
  6. 6.
    We inserted the created distractor into a random position in the context other than the beginning of the sample.
For example, When the sentence “James decides to run 3 sprints 3 times a week.” is selected from the sample in Table 3, a template “[name]  decides to run [num]  sprints [num]  times a week.” is crafted. Names and numbers are randomly selected from the candidates and placed into these placeholders, and the resulting distractor is then inserted into the context.
Overlap
To evaluate whether the model is influenced by the Overlap heuristic, we insert distractors $\tilde{p}$ into each sample following the steps below:
  1. 1.
    We substituted the placeholder [name]  within the Base distractor template with the person’s name found in the sample, appended by relational phrases such as “’s mother”, “’s father”, “’s son”, or “’s neighborhood” (e.g., in the sample from Table 3, this would become “James’s mother”).
  2. 2.
    We replaced the number in the sentence with another numerical value.
  3. 3.
    We placed the constructed distractor into context at the exact location where the Base distractor was positioned in the sample.
Position
To evaluate whether the model is induced by the heuristic of Position, we insert distractors $\tilde{p}$ into each sample. Each distractor is identical to the Base distractor except for its insertion point. Specifically, we relocated the insertion point of the distractor to a random position that is closer to the beginning of the context than the position used for the Base distractor.
Neg.
To evaluate the model’s response to the Neg. heuristic, we insert distractors $\tilde{p}$ into each sample created based on the following template:
[name]  doesn’t have [num] [object].
In this template:
  • [name]  is substituted with a random PN included in the sample.
  • [object]  is replaced with one of the following items: “apples,” “bananas,” “grapes,” “pencils,” or “books.”
  • [num]  is replaced with a different numerical value, using the same algorithm used for creating the Base distractor.
A.2 Evaluation
To determine whether the LMs selected the distractor during reasoning, we check if the numbers contained in the distractor $\tilde{p}$ are present in the facts $\bm{z}$. We calculate the percentage of samples where the distractor is selected.
Appendix B Experimental setting with artificial data (§ 4.1)
B.1 Dataset construction process
Base
We construct the artificial data by following the method outlined below, based on the template presented in Table 4.
  • Randomly assign one of the following names to the placeholders [nameA]  to [nameE]: “Alice,” “Bob,” “Carol,” “Dave,” “Eve,” “Frank,” “Grace,” “Heidi,” “Ivan,” “Judy,” “Kevin,” “Larry,” “Mallory,” “Nancy,” “Olivia,” “Peggy,” “Quentin,” “Rob,” “Sybil,” “Trent,” “Ursula,” “Victor,” “Walter,” “Xavier,” “Yvonne,” or “Zoe.”
  • Assign a randomly selected value from [nameA]  to [nameD]  to the placeholder [nameX].
  • Assign a random number from 0 to 100 to the placeholder [num].
  • Assign one of the objects “apples,” “bananas,” “grapes,” “pencils,” or “books” to the placeholder [object].
  • Assign either “more” or “less” to the placeholder [relation].
  • Randomly shuffle the order of the sentences.
Context:[nameA] has [num] [object].
[nameB] has [num] [relation] [object] than [nameA].
[nameC] has [num] [relation] [object] than [nameB].
[nameD] has [num] [relation] [object] than [nameC].
Question:How many [object] does [nameD] have?
distractor:[nameE] has [num] [relation] [object] than [nameX].
Table 4: Template of artificial data in § 4.1.
Overlap
We constructed a dataset to evaluate whether the model is induced by the Overlap heuristic by making certain modifications to the Base distractor for each sample. Specifically, we modified the value of [nameD]  by appending relational phrases such as “ ’s mother”, “ ’s father”, “ ’s son”, or “ ’s neighborhood” to the existing value of [nameD]. We then assigned this modified value to [nameE].
Position
To evaluate if the model is induced by the Position heuristic, we modify the Base distractor. Specifically, we altered the insertion point of the Base distractor to a randomly chosen position that is closer to the beginning of the context than the original position used in the Base distractor.
Neg.
To evaluate whether the model is induced by the Neg. heuristic, we construct a dataset by making modifications to the Base distractor. Specifically, we convert the Base distractor into a negative expression (e.g., [nameE]  doesn’t have [num]  [relation]  [object]  than [nameX]).
B.2 Evaluation
To determine whether the LMs selected the distractor during reasoning, we check if the subject of the distractor (i.e., [nameE]) is included in the facts $\bm{z}$. We calculate the percentage of samples where the distractor is selected.
Appendix C Artificial data in  § 4.2
Cntext:[nameA] has [num] [object].
[nameB] has [num] [relation] [object] than [nameA].
[nameC] has [num] [relation] [object] than [nameB].
[nameD] has [num] [relation] [object] than [nameC].
[nameE] has [num] [relation] [object] than [nameD].
Question:How many [object] does [nameE] have?
heurictic distractor:
[nameF] has [num] [relation] [object] than [nameA].
[nameG] has [num] [relation] [object] than [nameB].
[nameH] has [num] [relation] [object] than [nameC].
distractor:
[nameI] has [num] [relation] [object] than [nameD].
[nameJ] has [num] [relation] [object] than [nameF].
[nameK] has [num] [relation] [object] than [nameF].
[nameL] has [num] [relation] [object] than [nameG].
[nameM] has [num] [relation] [object] than [nameG].
[nameN] has [num] [relation] [object] than [nameJ].
[nameO] has [num] [relation] [object] than [nameJ].
[nameP] has [num] [relation] [object] than [nameK].
[nameQ] has [num] [relation] [object] than [nameK].
Table 5: Template of artificial data in § 4.2.
We prepare a template similar to a Table 5 and assign values to the template according to the following steps:
  • We create template as shown in table4.
  • Within the template, placeholders [nameA]  to [nameQ]  is filled randomly with names such as “Alice”, “Bob”, “Carol”, “Dave”, “Eve”, “Frank”, “Grace”, “Heidi”, “Ivan”, “Judy”, “Kevin”, “Larry”, “Mallory”, “Nancy”, “Olivia”, “Peggy”, “Quentin”, “Rob”, “Sybil”, “Trent”, “Ursula”, "Victor”, “Walter”, “Xavier”, “Yvonne”, “Zoe”.
  • The placeholder [num]  is filled with a random number from 0 to 100.
  • The placeholder [object]  is filled randomly with items such as “apples”, “bananas”, “grapes”, “pencils”, “books”.
  • The placeholder [relation]  is assigned either “more” or “less”.
  • Sentences within the context are shuffled randomly.
  • A distractor is inserted at a random position.
Then, using the following procedures, We create each expanded dataset. The heuristic distractors designed in this study are strongly influenced by each targeted heuristic. Each dataset consists of 300 problems.
  • For the Overlap dataset, the values “’s mother”, “’s father”, “’s son”, and “’s neighborhood” are appended to [nameE]  and assigned respectively to [nameF], [nameG], and [nameH]. Each of [nameF], [nameG], and [nameH]  hold different values.
  • For the Position dataset, the sentences with [nameF], [nameG], and [nameH]  as the subjects have distractors inserted closer to the beginning of the context than the sentences with [nameB], [nameC], and [nameD]  as the subjects. For other datasets, heuristic distractors are inserted at random positions.
  • For the Neg. dataset, the form of the heuristic distractor is changed to a negative form.
In § 4.2, the method to identify which premises are used for reasoning was similar to that in App. B, relying on regular expressions.
Appendix D Mearuing accuracy in § 4.1
In the experimental section, we not only measure the frequency with which $\tilde{p}$ is used in reasoning but also evaluate the accuracy. Table 6 below shows the accuracy rates of each model on GSM8K. Additionally, Table 7 below shows the accuracy rates of each model on artificial data.
From the Table 67, it is shown that GPT-4 had the highest accuracy rates across the datasets, while Llama2 had the lowest. It is expected that these outcomes are due to differences in the number of parameters in the model and the training data.
Base Over. Pos. Neg.
PaLM2 64.5% 59.2% 60.5% 71.1%
Llama2 30.3% 34.2% 30.3% 27.6%
GPT-3.5 81.6% 64.5% 81.6% 82.9%
GPT-4 85.5% 84.2% 82.9% 92.1%
Table 6: The accuracy while solving GSM8K.
Base Over. Pos. Neg.
PaLM2 60.0% 58.7% 77.0% 80.7%
Llama2 21.3% 14.3% 22.3% 20.0%
GPT-3.5 84.6% 87.3% 87.6% 82.9%
GPT-4 98.7% 94.0% 98.0% 99.7%
Table 7: The accuracy while solving artificial reasoning tasks.
Over.$\uparrow$ Pos.$\uparrow$ Neg.$\downarrow$
PaLM2 41.0% 14.0% 5.7%
Llama2 82.0% 93.3% 28.0%
GPT-3.5 11.7% 35.0% 0.0%
GPT-4 0.0% 0.0% 0.0%
Table 8: The frequency at which the model selected a distractor (i,e., $\tilde{p}$) while solving artificial reasoning tasks after changing few-shot examples.
Appendix E Generation settings
When using GPT-3.5, GPT-4, the settings are adjusted to temperature=0.0, frequency_penalty=0, and presence_penalty=0. Similarly, for PaLM2 and Llama2, the temperature is set to 0, with no sampling.
We use NVIDIA RTX A6000 (48GB) GPUs for inference with Llama2.
Appendix F Few-shot examples
The few-shot examples for models regarding datasets GSM8K and artificial data are shown in the respective Tables 910.
Answer the context question according to the following example.
Context: Leo’s assignment was divided into three parts. Weng earns \$12 an hour for babysitting. It took Leo twice as long to finish the second part. Yesterday, she just did 50 minutes of babysitting.
Question: How much did Weng earn?
Answer:
Weng earns 12/60 = 0.2 per minute.
Working 50 minutes, she earned 0.2 x 50 = 10.
The final answer is 10.
Context: Betty is saving money for a new wallet which costs \$100. Betty has only half of the money she needs. Alice is saving money for a new wallet which costs \$2000. Bettys parents decided to give her \$15 for that purpose, and her grandparents twice as much as her parents. Question: How much more money does Betty need to buy the wallet?
Question: How much more money does Betty need to buy the wallet?
Answer:
In the beginning, Betty has only 100 / 2 = 50.
Betty’s grandparents gave her 15 * 2 = 30.
This means, Betty needs 100 - 50 - 30 - 15 = 5 more.
The final answer is 5.
Context: Julie is reading a 120-page book. Yesterday, Julie was able to read 12 pages and today, she read twice as many pages as yesterday. Julie’ s mother makes \$18.00 an hour.
Question: How many pages are left to be read?
Answer:
Julie read 12 x 2 = <<12*2=24>>24 pages today
So she was able to read a total of 12 + 24 = 36 pages since yesterday.
There are 120 - 36 = 84 pages left to be read.
Since she wants to read half of the remaining pages tomorrow, then she should read 84/2 = 42 pages.
The final answer is 42.
Context: James writes a 2-page letter to 4 different friends, lived in America, twice a week. James writes a 3-page letter to 2 different friends, lived in Japan, twice a week.
Question: How many pages does James write each friend lived in Japan at one time?
Answer:
He writes each friend 3*2=6 pages a week.
So he writes 6*2=12 pages every week.
That means he writes 12*52=624 pages a year.
The final answer is 624.
Table 9: Examples of input given when solving GSM8K.
Answer the context question according to the following example.
Context: Walter has -22 apples. Ursula has 3 more apples than Walter. Victor has 3 more apples than Ursula. Quentin has 2 more apples than Ursula. Nancy has 3 more apples than Walter. Zoe has 3 more apples than Nancy. Heidi has 3 more apples than Nancy. Carol’s mother has 4 apples. Xavier has 3 more apples than Carol’s mother. Peggy has 4 more apples than Xavier. Dave has 13 more apples than Xavier. Bob has 1 more apples than Carol’s mother. Alice has 3 more apples than Bob. Sybil has 56 more apples than Bob.
Question: How many apples does Dave have?
Answer:
Carol’s mother has 4 apples, and Xavier has 3 more apples than Carol’s mother. So, Xavier has 4+3=7 apples.
Xavier has 7 apples, and Dave has 13 more apples than Xavier. So, Dave has 7+13=20 apples. The final answer is 20.
Context: Alice has 92 more bananas than Mallory. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 less bananas than Peggy. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry has 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Walter has 43 less bananas than Mallory. Nancy has 34 bananas. Grace has 41 more bananas than Xavier. Mallory has 55 less bananas than Nancy. Sybil has 3 less bananas than Nancy. Peggy has 50 more bananas than Walter. Trent has 33 less bananas than Xavier.
Question: How many bananas does Quentin have?
Answer:
Nancy has 34 bananas, and Mallory has 55 less bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 less bananas than Mallory. So, Walter has -21-43=-64 bananas.
Walter has -64 bananas, and Peggy has 50 more bananas than Walter. So, Peggy has -64+50=-14 bananas.
Peggy has -14 bananas, and Quentin has 35 less bananas than Peggy. So, Quentin has -14-35=-49 bananas.
The final answer is -49.
Context: Zoe has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 less apples than Zoe. Alice has 3 more apples than Grace. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier doesn’t have 3 more apples than Ivan.
Question: How many apples does Yvonne have?
Answer:
Eve has 2 apples, and Yvonne’s son has 3 more apples than Eve. So, Yvonne’s son has 2+3=5 apples.
Yvonne’s son has 5 apples, and Zoe has 10 more apples than Yvonne’s son. So, Zoe has 5+10=15 apples.
Zoe has 15 apples, and Yvonne has 3 less apples than Zoe. So, Yvonne has 15-3=12 apples.
The final answer is 12.
Context: Kevin’s friend has 33 less grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Sybil. Xavier has 36 more grapes than Heidi. Sybil has 55 less grapes than Trent. Kevin has 43 less grapes than Zoe. Heidi has 61 less grapes than Trent. Zoe has 88 more grapes than Sybil. Trent has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
Question: How many grapes does Kevin have?
Answer:
Rob has 55 grapes, and Trent has 40 more grapes than Rob. So, Trent has 55+40=95 grapes.
Trent has 95 grapes, and Sybil has 55 less grapes than Trent. So, Sybil has 95-55=40 grapes.
Sybil has 40 grapes, and Zoe has 88 more grapes than Sybil. So, Zoe has 40+88=128 grapes.
Zoe has 128 grapes, and Kevin has 43 less grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 128.
Table 10: Examples of input given when solving an artificial dataset.
Appendix G Effect of few-shot examples
We investigate whether the model’s heuristic is triggered by the few-shot examples. Specifically, we replace the few-shot examples in the following ways to study the relationship between the model’s heuristic and its inputs:
1. We change the few-shot examples to induce Overlap (as shown in Table 11) and examine whether this increases the reasoning frequency with the use of distractors in the Overlap dataset compared to what is shown in Table 1.
2. We change the few-shot examples to induce Position (as shown in Table 12) and check if there’s an increase in reasoning frequency with the use of distractors in the Position dataset compared to Table 1.
3. We change the few-shot examples to induce Neg. (as shown in Table 13), and investigate if there’s a decrease in reasoning frequency with the use of distractors in the Neg. dataset compared to Table 1.
We measure the frequency of selecting $\tilde{p}$ in the settings of  § 4.1. The results are presented in Table 8. As shown in Tables 18, although the few-shot examples fed into the models such as GPT-3.5, GPT-4, and PaLM2 was changed, there was no significant change in reasoning frequency as described. This suggests that the model’s heuristic does not merely mimic the examples provided as input. On the other hand, the Llama2 model was more prone to being misled by changes in input, and smaller models demonstrated a reduced capacity to reach the correct answers directory.
Answer the context question according to the following example.
Context: Context: Walter has -22 apples. Ursula has 3 more apples than Walter. Victor has 3 more apples than Ursula. Quentin has 2 more apples than Ursula. Nancy has 3 more apples than Walter. Zoe has 3 more apples than Nancy. Heidi has 3 more apples than Nancy. Dave’s mother has 4 apples. Dave’s father has 3 more apples than Dave’s mother. Peggy has 4 more apples than Dave’s father. Dave has 13 more apples than Dave’s father. Bob has 1 more apples than Carol’s mother. Alice has 3 more apples than Bob. Sybil has 56 more apples than Bob.
Question: How many apples does Dave have?
Answer:
Dave’s mother has 4 apples, and Dave’s father has 3 more apples than Dave’s mother. So, Dave’s father has 4+3=7 apples.
Dave’s father has 7 apples, and Dave has 13 more apples than Dave’s father. So, Dave has 7+13=20 apples.
The final answer is 10.
Context: Alice has 92 more bananas than Quentin’s mother. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 less bananas than Quentin’s father. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Quentin’s father. Larry has 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Walter has 43 less bananas than Quentin’s mother. Nancy has 34 bananas. Grace has 41 more bananas than Xavier. Quentin’s mother has 55 less bananas than Nancy. Sybil has 3 less bananas than Nancy. Quentin’s father has 50 more bananas than Walter. Trent has 33 less bananas than Xavier.
Question: How many bananas does Quentin have?
Answer:
Nancy has 34 bananas, and Quentin’s mother has 55 less bananas than Nancy. So, Quentin’s mother has 34-55=-21 bananas.
Quentin’s mother has -21 bananas, and Walter has 43 less bananas than Quentin’s mother. So, Walter has -21-43=-64 bananas.
Walter has -64 bananas, and Quentin’s father has 50 more bananas than Walter. So, Quentin’s father has -64+50=-14 bananas.
Quentin’s father has -14 bananas, and Quentin has 35 less bananas than Quentin’s father. So, Quentin has -14-35=-49 bananas.
The final answer is -49.
Context: Yvonne’s father has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 less apples than Yvonne’s father. Alice has 3 more apples than Grace. Trent has 34 more apples than Yvonne’s father. Ivan has 3 apples. Ursula has 3 more apples than Yvonne’s father. Grace has 3 apples. Xavier has 3 more apples than Ivan.
Question: How many apples does Yvonne have?
Answer:
Eve has 2 apples, and Yvonne’s son has 3 more apples than Eve. So, Yvonne’s son has 2+3=5 apples.
Yvonne’s son has 5 apples, and Yvonne’s father has 10 more apples than Yvonne’s son. So, Yvonne’s father has 5+10=15 apples.
Yvonne’s father has 15 apples, and Yvonne has 3 less apples than Yvonne’s father. So, Yvonne has 15-3=12 apples.
The final answer is 12.
Context: Kevin’s friend has 33 less grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Kevin’s neighborhood. Xavier has 36 more grapes than Heidi. Kevin’s neighborhood has 55 less grapes than Kevin’s friend. Kevin has 43 less grapes than Kevin’s mother. Heidi has 61 less grapes than Kevin’s friend. Kevin’s mother has 88 more grapes than Kevin’s neighborhood. Kevin’s friend has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
Question: How many grapes does Kevin have?
Answer:
Rob has 55 grapes, and Kevin’s friend has 40 more grapes than Rob. So, Kevin’s friend has 55+40=95 grapes.
Kevin’s friend has 95 grapes, and Kevin’s neighborhood has 55 less grapes than Kevin’s friend. So, Kevin’s neighborhood has 95-55=40 grapes.
Kevin’s neighborhood has 40 grapes, and Kevin’s mother has 88 more grapes than Kevin’s neighborhood. So, Kevin’s mother has 40+88=128 grapes.
Kevin’s mother has 128 grapes, and Kevin has 43 less grapes than Kevin’s mother. So, Kevin has 128-43=85 grapes.
The final answer is 128.
Table 11: Examples of input given when solving the Overlap dataset.
Answer the context question according to the following example.
Context: Carol’s mother has 4 apples. Xavier has 3 more apples than Carol’s mother. Dave has 13 more apples than Xavier. Walter has -22 apples. Ursula has 3 more apples than Walter. Victor has 3 more apples than Ursula. Quentin has 2 more apples than Ursula. Nancy has 3 more apples than Walter. Zoe has 3 more apples than Nancy. Heidi has 3 more apples than Nancy. Peggy has 4 more apples than Xavier. Bob has 1 more apples than Carol’s mother. Alice has 3 more apples than Bob. Sybil has 56 more apples than Bob.
Question: How many apples does Dave have?
Answer:
Carol’s mother has 4 apples, and Xavier has 3 more apples than Carol’s mother. So, Xavier has 4+3=7 apples.
Xavier has 7 apples, and Dave has 13 more apples than Xavier. So, Dave has 7+13=20 apples. The final answer is 20.
Context: Nancy has 34 bananas. Mallory has 55 less bananas than Nancy. Walter has 43 less bananas than Mallory. Peggy has 50 more bananas than Walter. Quentin has 35 less bananas than Peggy. Alice has 92 more bananas than Mallory. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne has 79 more bananas than Sybil. Judy has 23 more bananas than Alice. Dave has 60 more bananas than Victor. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry has 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Grace has 41 more bananas than Xavier. Sybil has 3 less bananas than Nancy. Trent has 33 less bananas than Xavier.
Question: How many bananas does Quentin have?
Answer:
Nancy has 34 bananas, and Mallory has 55 less bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 less bananas than Mallory. So, Walter has -21-43=-64 bananas.
Walter has -64 bananas, and Peggy has 50 more bananas than Walter. So, Peggy has -64+50=-14 bananas.
Peggy has -14 bananas, and Quentin has 35 less bananas than Peggy. So, Quentin has -14-35=-49 bananas.
The final answer is -49.
Context: Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Zoe has 10 more apples than Yvonne’s son. Yvonne has 3 less apples than Zoe. Alice has 3 more apples than Grace. Quentin has 3 more apples than Yvonne. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier has 3 more apples than Ivan.
Question: How many apples does Yvonne have?
Answer:
Eve has 2 apples, and Yvonne’s son has 3 more apples than Eve. So, Yvonne’s son has 2+3=5 apples.
Yvonne’s son has 5 apples, and Zoe has 10 more apples than Yvonne’s son. So, Zoe has 5+10=15 apples.
Zoe has 15 apples, and Yvonne has 3 less apples than Zoe. So, Yvonne has 15-3=12 apples.
The final answer is 12.
Context: Rob has 55 grapes. Trent has 40 more grapes than Rob. Sybil has 55 less grapes than Trent. Zoe has 88 more grapes than Sybil. Kevin has 43 less grapes than Zoe. Kevin’s friend has 33 less grapes than Rob. Ivan has 43 more grapes than Victor. Victor has 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia has 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy has 78 more grapes than Olivia. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Sybil. Xavier has 36 more grapes than Heidi. Heidi has 61 less grapes than Trent. Walter has 38 more grapes than Victor.
Question: How many grapes does Kevin have?
Answer:
Rob has 55 grapes, and Trent has 40 more grapes than Rob. So, Trent has 55+40=95 grapes.
Trent has 95 grapes, and Sybil has 55 less grapes than Trent. So, Sybil has 95-55=40 grapes.
Sybil has 40 grapes, and Zoe has 88 more grapes than Sybil. So, Zoe has 40+88=128 grapes.
Zoe has 128 grapes, and Kevin has 43 less grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 128.
Table 12: Examples of input given when solving the Position dataset.
Answer the context question according to the following example.
Context: Walter doesn’t have -22 apples. Ursula has 3 more apples than Walter. Victor has 3 more apples than Ursula. Quentin has 2 more apples than Ursula. Nancy doesn’t have 3 more apples than Walter. Zoe has 3 more apples than Nancy. Heidi doesn’t have 3 more apples than Nancy. Carol’s mother has 4 apples. Xavier has 3 more apples than Carol’s mother. Peggy has 4 more apples than Xavier. Dave has 13 more apples than Xavier. Bob doesn’t have 1 more apples than Carol’s mother. Alice has 3 more apples than Bob. Sybil has 56 more apples than Bob.
Question: How many apples does Dave have?
Answer:
Carol’s mother has 4 apples, and Xavier has 3 more apples than Carol’s mother. So, Xavier has 4+3=7 apples.
Xavier has 7 apples, and Dave has 13 more apples than Xavier. So, Dave has 7+13=20 apples. The final answer is 20.
Context: Alice has 92 more bananas than Mallory. Victor has 10 less bananas than Walter. Xavier has 59 more bananas than Sybil. Yvonne doesn’t have 79 more bananas than Sybil. Judy doesn’t have 23 more bananas than Alice. Dave has 60 more bananas than Victor. Quentin has 35 less bananas than Peggy. Heidi has 95 more bananas than Victor. Ursula doesn’t have 32 more bananas than Peggy. Larry doesn’t have 17 less bananas than Alice. Zoe has 58 less bananas than Yvonne. Ivan has 43 less bananas than Yvonne. Walter has 43 less bananas than Mallory. Nancy has 34 bananas. Grace doesn’t have 41 more bananas than Xavier. Mallory has 55 less bananas than Nancy. Sybil doesn’t have 3 less bananas than Nancy. Peggy has 50 more bananas than Walter. Trent doesn’t have 33 less bananas than Xavier.
Question: How many bananas does Quentin have?
Answer:
Nancy has 34 bananas, and Mallory has 55 less bananas than Nancy. So, Mallory has 34-55=-21 bananas.
Mallory has -21 bananas, and Walter has 43 less bananas than Mallory. So, Walter has -21-43=-64 bananas.
Walter has -64 bananas, and Peggy has 50 more bananas than Walter. So, Peggy has -64+50=-14 bananas.
Peggy has -14 bananas, and Quentin has 35 less bananas than Peggy. So, Quentin has -14-35=-49 bananas.
The final answer is -49.
Context: Zoe has 10 more apples than Yvonne’s son. Eve has 2 apples. Yvonne’s son has 3 more apples than Eve. Quentin has 3 more apples than Yvonne. Yvonne has 3 less apples than Zoe. Alice has 3 more apples than Grace. Trent has 34 more apples than Zoe. Ivan has 3 apples. Ursula has 3 more apples than Zoe. Grace has 3 apples. Xavier doesn’t have 3 more apples than Ivan.
Question: How many apples does Yvonne have?
Answer:
Eve has 2 apples, and Yvonne’s son has 3 more apples than Eve. So, Yvonne’s son has 2+3=5 apples.
Yvonne’s son has 5 apples, and Zoe has 10 more apples than Yvonne’s son. So, Zoe has 5+10=15 apples.
Zoe has 15 apples, and Yvonne has 3 less apples than Zoe. So, Yvonne has 15-3=12 apples.
The final answer is 12.
Context: Kevin’s friend has 33 less grapes than Rob. Ivan doesn’t have 43 more grapes than Victor. Victor doesn’t have 33 less grapes than Kevin’s friend. Ursula has 75 less grapes than Zoe. Alice has 11 more grapes than Eve. Dave has 11 more grapes than Eve. Olivia doesn’t have 29 more grapes than Kevin’s friend. Mallory has 97 more grapes than Olivia. Judy doesn’t have 78 more grapes than Olivia. Rob has 55 grapes. Frank has 70 less grapes than Heidi. Eve has 84 less grapes than Sybil. Xavier doesn’t 36 more grapes than Heidi. Sybil has 55 less grapes than Trent. Kevin has 43 less grapes than Zoe. Heidi has 61 less grapes than Trent. Zoe has 88 more grapes than Sybil. Trent has 40 more grapes than Rob. Walter has 38 more grapes than Victor.
Question: How many grapes does Kevin have?
Answer:
Rob has 55 grapes, and Trent has 40 more grapes than Rob. So, Trent has 55+40=95 grapes.
Trent has 95 grapes, and Sybil has 55 less grapes than Trent. So, Sybil has 95-55=40 grapes.
Sybil has 40 grapes, and Zoe has 88 more grapes than Sybil. So, Zoe has 40+88=128 grapes.
Zoe has 128 grapes, and Kevin has 43 less grapes than Zoe. So, Kevin has 128-43=85 grapes.
The final answer is 128.
Table 13: Examples of input given when solving the Neg dataset.